OpenStudy (anonymous):

Find the quadratic equation of roots 3 and 4, -2 and 7, -9 and 1.

OpenStudy (ajprincess):

3,4 (x-3)(x-4)=0 x^2-3x-4x+12=0 x^2-7x+12=0 If a, b are roots of a quadratic eqn (x-a)(x-b)=0 x^2-ax-bx+ab=0 x^2-(a+b)x+ab=0 See if u cn do the rest two.

OpenStudy (anonymous):

Thanks

OpenStudy (ajprincess):

yw